to resistors two resistors R1 equals 20 ohms and R2 equals 20 ohms are connected in series with an EMF of 3.0 volts what is the equivalent resistor in the circuit
Answer #1

When two resistors are connected in series, then equivalent resistance is given by:

Req = R1 + R2

Req = 20 + 20

Req = 40 ohm (your answer)

Let me know if you've any query.

Extra information: If both resistor are in parallel, then equivalent resistance will be:

1/Req = 1/R1 + 1/R2

1/Req = 1/20 + 1/20 = 2/20 = 1/10

Req = 10 ohm